# What Impulse Response Do Instrumental Variables Identify?

### v3

@@
Bonsoo Koo Seojeong Lee
Myung Hwan Seo Masaya Takano
Department of Econometrics and Business Statistics, Monash University, bonsoo.koo@monash.eduDepartment of Economics, Seoul National University, s.jay.lee@snu.ac.krDepartment of Economics, Seoul National University, myunghseo@snu.ac.krDepartment of Economics, Seoul National University, mtakano@snu.ac.kr
(July 29, 2025111We acknowledge that this research is supported under the Australian Research Council Discovery Project funding scheme (project number DP210101440) and by the Seoul National University Research Grant in 2021. )
Macroeconomic shocks are often composites of multiple components. We show that the local projection-IV (LP-IV) estimand aggregates component-wise impulse responses with potentially negative weights, challenging its causal interpretation. To address this, we propose identification strategies using multiple sign-restricted IVs or disaggregated data, which recover structurally meaningful responses even when individual LP-IV estimands are non-causal. We also show that, under weak stationarity, the identified sets are sharp and cannot be further narrowed in some key cases. Applications to fiscal and monetary policy demonstrate the practical value of our approach.
@@ 1  Introduction
Nevertheless, much of the LP-IV literature—including key contributions by Stock and Watson (2018) and Plagborg-Møller and Wolf (2021, 2022)—continues to operate under the assumption that the shocks identified through IVs are homogeneous. This overlooks the composite nature of many macroeconomic shocks, leading to potential biases in the LP-IV estimands. Recognizing this limitation, Miranda-Agrippino and Ricco (2021) have highlighted the potential bias in LP-IV estimates due to this composite nature and proposed modifications to the IVs themselves, utilizing more information to better align the instruments with the exclusion restrictions required for valid identification. Miranda-Agrippino and Ricco (2023) further emphasized that the IVs developed in their 2021 paper could identify the shock of interest when some shocks are invertible provided that the shock of interest can be recovered from reduced-form shocks. Instead, we do not assume such a partial invertibility and allow for set-identified impulse responses using arguably less restrictive sign restrictions—acknowledging that IVs may be correlated with multiple shocks, with only the signs of these correlations known.
We propose a more precise interpretation of the LP-IV estimand by employing a potentially non-invertible vector moving average (SVMA) model. This model allows for the IVs to be correlated with multiple component shocks within the policy variable, reflecting the composite nature of macroeconomic shocks. We demonstrate that the impulse response identified by the LP-IV method can be expressed as a weighted sum of impulse responses to the individual components of the composite structural shock. These weights are determined by the correlation between the IV and each component shock. Importantly, since this correlation is not restricted to be non-negative, the LP-IV estimand does not necessarily identify an informative quantity like a proper weighted average of the impulse responses we aim to recover.
@@
$\begin{array}[]{ccccccccc}\widehat{\beta}_{h}&=&\widehat{w}_{\text{defense}}&\times&\widehat{\theta}_{h,\text{defense}}&+&\widehat{w}_{\text{non-defense}}&\times&\widehat{\theta}_{h,\text{non-defense}}.\\ 0.37&&1.87&&0.68&&-0.87&&1.02\\ \end{array}$ (3)
@@
Our finding that only LP-IV estimands with non-negative weights have a structural interpretation highlights the need for careful instrument selection. When a structural shock comprises multiple components, the sign of the correlation between each component and a given instrument determines whether the resulting estimate is causally meaningful. As such, researchers must go beyond traditional IV validity and examine the structure of the shock-instrument relationship.
@@
The first strategy is to use disaggregated data to identify the weights. For example, quarterly data on defense and non-defense spending in the US after the WWII are available and we can use these variables to identify the weights for the defense and non-defense spending multipliers. When the number of available IVs is at least as many as the number of components in the structural shock, the component-wise impulse responses can be identified regardless of the signs of the weights.
@@
Our second strategy uses sign restrictions to derive identified sets for component-wise impulse responses. While sign restrictions are commonly used in the SVAR literature (e.g., Uhlig, 2005) and have been discussed by Plagborg-Møller and Wolf (2021) and Alpanda, Granziera, and Zubairy (2021) in the LP framework, our approach differs by imposing sign restrictions on the correlations between shocks and IVs rather than directly on the structural impulse responses. These restrictions are coherent with the IV-based identification and easier to justify. Similar types of sign restrictions are used in work focusing on Bayesian SVARs, such as Jarociński and Karadi (2020), Arias, Rubio-Ramírez, and Waggoner (2021), and Giacomini, Kitagawa, and Read (2022). Additionally, our strategy requires no additional computation since the LP-IV estimates for each IV are directly used as bound estimates, unlike Bayesian approaches that often necessitate extensive computation or simulation. Surprisingly, an IV that is correlated with the component shocks with opposing signs can be more useful than one correlated with the component shocks in the same direction. When two IVs with opposing signs are combined, the resulting identified set is the most informative, as shown in Proposition 2. Lastly, our approach easily accommodates sign restrictions with many IVs, as a tighter identified set can be obtained by intersecting the bounds.
@@
In addition to proposing new identification strategies, we assess whether the LP-IV based identified sets can be further sharpened under weak stationarity. Focusing on a finite-order SVMA model with two IVs and two component shocks, and imposing sign restrictions as in Jarociński and Karadi (2020), we show that the LP-IV identified set for the component-wise impulse responses takes the form of a half-rectangle. We demonstrate that the identified set is inevitably unbounded under certain sufficient conditions. Separately, we provide alternative sufficient conditions under which the identified set for the other component-wise impulse response is sharp. This unboundedness result is empirically relevant and has not been recognized in the existing literature.
As argued by Wolf (2022), the standard Bayesian procedure employed in such analyses often fails to adequately address the fact that sign restrictions lead to partial or set identification of the impulse response rather than point identification. Giacomini, Kitagawa, and Reed (2022) propose a method to robustify the Bayesian approach to better handle this set identification issue. However, a significant challenge remains: the identified set is not always bounded, which violates a key regularity condition required for the validity of their proposed confidence region.
@@
We focus on LP-IV due to its simplicity in estimation and structural interpretation, as well as its capacity to combine information from multiple IVs, enabling the identification of structural parameters even when individual IVs do not provide useful identification. Moreover, it is easier to allow for nonlinear functional forms in the LP specification. This renders the LP-IV estimators more robust to misspecification of the data generating process (Nakamura and Steinsson, 2018).
@@ 2.1 Structural Vector Moving Average
[…]
@@ 2.2 Local Projections with an Instrumental Variable
    $E[z_{t}\varepsilon_{s,t}]=0$ for all $s=S+1,S+2,...,m$ (contemporaneous exogeneity)
@@
Define $\alpha_{s}=E[z_{t}\varepsilon_{s,t}]$. Assumption 1(i) implies that $\alpha_{s}\neq 0$ for some $s=1,2,...,S$. Assumption 1 is an extension of Condition LP-IV of Stock and Watson (2018) allowing that more than one structural shocks are correlated with the IV and the correlations are heterogeneous across the shocks.
@@
If random variables $y_{t}$ and $x_{t}$ are elements of $\bm{Y}_{t}$ generated according to (5) and a random variable $z_{t}$ satisfies Assumption 1, and $\theta_{0,xs}>0$ for $s=1,2,...,S$, then for $h=0,1,2,...$,
@@
The SVMA model offers a structural interpretation of the impulse responses $\theta_{h,ys}$ in Proposition 1. However, a similar result to (6) can be obtained even without relying on a structural model. [^5] We thank an anonymous referee for this insightful comment. Define $\theta_{h,ys}$ as the coefficient from the orthogonal projection, given by $\text{cov}(y_{t+h},\varepsilon_{s,t})/\text{var}(\varepsilon_{s,t})$. Suppose $z_{t}$ is a linear combination of the component shocks: $z_{t}=\lambda’\bm{e}_{t}$, where $\bm{e}_{t}\equiv(\varepsilon_{1,t},\varepsilon_{2,t},\cdots,\varepsilon_{S,t})’$ ’. Since the shocks $\varepsilon_{s,t}$ are mutually uncorrelated, the orthogonal projection of $y_{t+h}$ on $\bm{e}_{t}$ is $\mathcal{P}(y_{t+h}|\bm{e}_{t})=\sum_{s=1}^{S}\theta_{h,ys}\varepsilon_{s,t}$. The projection error $y_{t+h}-\mathcal{P}(y_{t+h}|\bm{e}_{t})$ is orthogonal to the subspace spanned by $\bm{e}_{t}$, and hence also to $z_{t}$. Therefore, the projection of $y_{t+h}$ onto $z_{t}$ is:
@@
For all $s=1,2,...,S$, either $\alpha_{s}\theta_{0,xs}\geq 0$ or $\alpha_{s}\theta_{0,xs}\leq 0$.
@@ 3.1 It is NOT okay to use 2SLS
$E[z_{t}^{A}\varepsilon_{1,t}]=$ $0.5,~~E[z_{t}^{A}\varepsilon_{2,t}]=0.5,~~E[z_{t}^{B}\varepsilon_{1,t}]=0.125,~~E[z_{t}^{B}\varepsilon_{2,t}]=1.5$
@@
In summary, it is generally not advisable to combine multiple IVs using conventional estimators such as 2SLS. Instead, there are more constructive approaches to leverage the information provided by multiple IVs. One possibility is to explore additional sources of information that can assist in identifying the impulse responses to the components of the structural shock. In the remainder of this section, we propose two such strategies: (i) using disaggregated data and restrictions on the inter-component impulse responses, and (ii) imposing sign restrictions on the correlation between the IV and the component-wise shocks. For simplicity, our identification strategies focus on the case where $S=2$ (two components), which covers the main applications discussed in Sections 4 and 5.888The strategies developed in this section can be extended to the cases with $S\geq 3$ in principle, but with considerably more complex notations.
@@ 3.2 Identification by Disaggregated Data
[…]
@@
for $s=1,2$ and $r\neq s$, provided that $\psi_{0,21}\psi_{0,12}\neq 1$. By imposing restrictions on $\psi_{0,sr}$, $\alpha_{s}$ are identified, and thus, $\theta_{h,y1}$ and $\theta_{h,y2}$ in (16) are identified.
@@ 3.3 Identification Bounds by Sign Restrictions
Panels (d)-(i) show the identified sets when the different sign restrictions, varying across the instruments, are imposed. In these cases multiple IV’s provide more informative bound for the component-wise impulse responses without the inequality relationship between $\theta_{h,y1}/\theta_{0,x1}$ and $\theta_{h,y2}/\theta_{0,x2}$. Furthermore, the intersection of two identified sets with opposing sign restrictions gives the identified set bounded both above and below. This can happen when $\Theta_{+-}^{A}$ and $\Theta_{-+}^{B}$ intersect. If $\beta_{h}^{A}>\beta_{h}^{B}$, then the identified set for $(\theta_{h,y1}/\theta_{0,x1},\theta_{h,y2}/\theta_{0,x2})$ is given by
@@
Suppose that random variables $z_{t}^{A}$ and $z_{t}^{B}$ satisfy Assumption 1. We assume that $\text{cov}(z_{t}^{A},x_{t})>0$, $\text{cov}(z_{t}^{B},x_{t})>0$, and $\beta_{h}^{A}>\beta_{h}^{B}$ without loss of generality. [^9] If the covariance is not positive, we use $-z_{t}$ as an instrument. Note that this will change the sign of $\alpha_{1}$ and $\alpha_{2}$ but not the sign of $\beta_{h}$.
@@ 3.4 Sharpness of Sign-Restricted LP-IV Identification
While we state sharp identification results for a bivariate VMA(1)(1)( 1 ) model, this is solely to simplify the exposition; our results hold for a more general VMA model, as discussed in the Supplemental Appendix.
@@ 3.4.1 Model Setup
Consider the SVMA(1)(1)( 1 ) model for the pair $(\mathbf{Y}_{t}^{\prime},\mathbf{Z}_{t}^{\prime})^{\prime}$ of 2×12$\times$ 12 × 1 endogenous and instrumental variable vectors $\mathbf{Y}_{t}=(x_{t},y_{t})^{\prime}$ and $\mathbf{Z}_{t}=(z_{t}^{A},z_{t}^{B})^{\prime}$ with respect to the 3×13$\times$ 13 × 1 structural shocks $\bm{\varepsilon}_{t}$ and the 2×12$\times$ 12 × 1 projection error vector $\mathbf{v}_{t}$:
@@
Note that this specification imposes the exclusion restriction that $\alpha_{3}^{j}=0$ for $j=A,B$, and $\alpha_{s}^{j}=d_{s}^{j}\lambda_{s}$ for $j=A,B$ and $s=1,2$. The parameters of interest are $\theta_{1,y1}/\theta_{0,x1}$ and $\theta_{1,y2}/\theta_{0,x2}$, the relative impulse responses of $y_{t}$ with respect to $\varepsilon_{1,t}$ and $\varepsilon_{2,t}$ at horizon 111. We may consider the impulse responses at horizon 0 instead, and the results in Section 3.4.2 remain valid.
@@
$\begin{cases}(\beta^{A},\beta^{B})\times(-\infty,\beta^{A})&\text{if }\beta^{A}<\beta^{B}\\ (\beta^{B},\beta^{A})\times(\beta^{A},\infty)&\text{if }\beta^{A}>\beta^{B}\end{cases}$ (20)

### v4

@@
Bonsoo Koo Seojeong Lee Myung Hwan Seo Masaya Takano Department of Econometrics and Business Statistics, Monash University, bonsoo.koo@monash.eduDepartment of Economics, Seoul National University, s.jay.lee@snu.ac.krDepartment of Economics, Hong Kong University of Science and Technology, myunghseo@ust.hkDepartment of Economics, Seoul National University, mtakano@snu.ac.kr
The local projection–instrumental variable (LP-IV) literature has been largely silent on cases in which impulse responses are set-identified, arising when the shock of interest is composite and instruments are correlated with multiple components. We demonstrate that LP-IV estimands constructed using one instrument at a time identify affine combinations of impulse responses to structural shock components with instrument-specific and potentially negative weights, challenging standard causal interpretation. The two-stage least squares compounds the identification problem. However, we show that individual LP-IV estimands characterize the identified set when sign restrictions on the correlations between instruments and structural shock components are imposed. Under weak stationarity, these identified sets are sharp and cannot be further narrowed in key cases. Two empirical examples—decomposing the U.S. government spending multiplier and disentangling pure monetary shocks from central bank information shocks—illustrate the usefulness of our approach.
@@ 1  Introduction
Nevertheless, much of the LP-IV literature including key contributions by Stock and Watson (2018) and Plagborg-Møller and Wolf (2021, 2022) continues to operate under the assumption that the shocks identified through IVs are homogeneous. This significantly constrains the applicability of the LP-IV approach. Recognizing this limitation, Miranda-Agrippino and Ricco (2021) have highlighted the potential bias in LP-IV estimates due to the composite nature and proposed modifications to the IVs themselves, utilizing more information to better align the instruments with the exclusion restrictions required for valid identification. Miranda-Agrippino and Ricco (2023) further emphasized that the IVs developed in their 2021 paper could identify the shock of interest when some shocks are invertible provided that the shock of interest can be recovered from reduced-form shocks.
We start by deriving an explicit representation of the LP-IV estimand by employing a potentially non-invertible vector moving average (SVMA) model. This model allows for the IVs to be correlated with multiple component shocks within the policy variable, reflecting the composite nature of macroeconomic shocks. We show that the LP-IV estimand is expressed as a weighted sum of impulse responses to the individual components of the composite structural shock. These weights are determined by the correlation between the IV and each component shock. Importantly, since this correlation is not restricted to be non-negative, the LP-IV estimand does not necessarily identify an informative quantity like a proper weighted average of the impulse responses.
@@
[…]
@@
The first strategy is to use disaggregated data to identify the weights. For example, quarterly data on defense and non-defense spending in the US after the WWII are available and we can use these variables to identify the weights for the defense and non-defense spending multipliers. When the number of available IVs is at least as many as the number of components in the structural shock, the impulse responses to each component shock can be identified regardless of the signs of the weights.
@@
Our second strategy uses sign restrictions, which are commonly used in the SVAR literature (e.g., Uhlig, 2005) and have been discussed by Plagborg-Møller and Wolf (2021) and Alpanda, Granziera, and Zubairy (2021) in the LP framework. Specifically, our approach builds on sign restrictions on the correlations between shocks and IVs rather than directly on the structural impulse responses. These restrictions are coherent with the IV-based identification and easy to justify. This types of sign restrictions are used in work focusing on Bayesian SVARs, such as Jarociński and Karadi (2020), Arias, Rubio-Ramírez, and Waggoner (2021), and Giacomini, Kitagawa, and Read (2022).
Moreover, our strategy requires no additional computation since the LP-IV estimates for each IV are directly used as bound estimates, unlike Bayesian approaches that often necessitate extensive computation or simulation. Surprisingly, an IV that is correlated with the component shocks with opposing signs can be more useful than one correlated with the component shocks in the same direction. When two IVs with opposing signs are combined, the resulting identified set is the most informative, as shown in Proposition 2. Lastly, our approach easily accommodates sign restrictions with many IVs, as a tighter identified set can be obtained by intersecting the bounds.
@@
In addition to proposing new identification strategies, we assess whether the LP-IV based identified sets can be further sharpened under weak stationarity. Focusing on a finite-order SVMA model with two IVs and two component shocks, and imposing sign restrictions as in Jarociński and Karadi (2020), we show that the LP-IV identified set for the component-wise impulse responses takes the form of a half-rectangle. We demonstrate that the identified set is inevitably unbounded under certain sufficient conditions. Separately, we provide alternative sufficient conditions under which the identified set for the other component-wise impulse response is sharp. This unboundedness result is empirically relevant and has not been fully recognized in the existing literature.
As argued by Wolf (2022), the standard Bayesian procedure employed in such analyses often fails to adequately address the fact that sign restrictions lead to partial or set identification of the impulse response rather than point identification. Giacomini, Kitagawa, and Read (2022) propose a method to robustify the Bayesian approach to better handle this set identification issue. However, a significant challenge remains: the identified set is not always bounded, which violates a key regularity condition required for the validity of their proposed confidence region. The possibility of unbounded identified sets is implicitly suggested by Baumeister and Hamilton (2015) and is analyzed explicitly by Read (2024) who also demonstrates the empirical relevance of unboundedness in set-identified SVARs when impulse responses are normalised to a unit effect.
@@ 2.1 Structural Vector Moving Average
Here we are slightly abusing notation by letting $\theta_{h,ys}$ be the $(n,s)$-th element of $\bm{\Theta}_{h}$. This is to emphasize that $y_{t}$ is the main variable of interest.
@@ 2.2 Local Projections with an Instrumental Variable
    $E[z_{t}\varepsilon_{s,t}]=0$ for all s=S+1,S+2,,ms=S+1,S+2,...,m (contemporaneous exogeneity)
@@
Define $\alpha_{s}=E[z_{t}\varepsilon_{s,t}]$. Assumption 1(i) implies that $\alpha_{s}\neq 0$ for some s=1,2,,Ss=1,2,...,S. Assumption 1 is an extension of Condition LP-IV of Stock and Watson (2018) allowing that more than one structural shocks are correlated with the IV and the correlations are heterogeneous across the shocks.
@@
If random variables $y_{t}$ and $x_{t}$ are elements of $\bm{Y}_{t}$ generated according to (5) and a random variable $z_{t}$ satisfies Assumption 1, and $\theta_{0,xs}>0$ for s=1,2,,Ss=1,2,...,S, then for $h=0,1,2,$...,
@@
The SVMA model offers a structural interpretation of the impulse responses $\theta_{h,ys}$ in Proposition 1. However, a similar result to (6) can be obtained even without relying on a structural model. [^4] We thank an anonymous referee for this insightful comment. Define $\theta_{h,ys}$ as the coefficient from the orthogonal projection, given by $\text{cov}(y_{t+h},\varepsilon_{s,t})/\text{var}(\varepsilon_{s,t})$. Suppose $z_{t}$ is a linear combination of the component shocks: $z_{t}=\lambda’\bm{e}_{t}$, where $\bm{e}_{t}\equiv(\varepsilon_{1,t},\varepsilon_{2,t},\cdots,\varepsilon_{S,t})$’. Since the shocks $\varepsilon_{s,t}$ are mutually uncorrelated, the orthogonal projection of $y_{t+h}$ on $\bm{e}_{t}$ is $\mathcal{P}(y_{t+h}|\bm{e}_{t})=\sum_{s=1}^{S}\theta_{h,ys}\varepsilon_{s,t}$. The projection error $y_{t+h}-\mathcal{P}(y_{t+h}|\bm{e}_{t})$ is orthogonal to the subspace spanned by $\bm{e}_{t}$, and hence also to $z_{t}$. Therefore, the projection of $y_{t+h}$ onto $z_{t}$ is:
@@
For all s=1,2,,Ss=1,2,...,S, either $\alpha_{s}\theta_{0,xs}\geq 0$ or $\alpha_{s}\theta_{0,xs}\leq 0$.
@@ 3.1 It is NOT okay to use 2SLS
$E[z_{t}^{A}\varepsilon_{1,t}]=$ $0.5,~~E[z_{t}^{A}\varepsilon_{2,t}]=0.5,~~E[z_{t}^{B}\varepsilon_{1,t}]=0.125,~~E[z_{t}^{B}\varepsilon_{2,t}]=1.5.$
@@
In summary, it is generally not advisable to combine multiple IVs using conventional estimators such as 2SLS. Instead, there are more constructive approaches to leverage the information provided by multiple IVs. One possibility is to explore additional sources of information that can assist in identifying the impulse responses to the components of the structural shock. In the remainder of this section, we propose two such strategies: (i) using disaggregated data and restrictions on the inter-component impulse responses, and (ii) imposing sign restrictions on the correlation between the IV and the component-wise shocks. For simplicity, our identification strategies focus on the case where $S=2$ (two components), which covers the main applications discussed in Sections 4 and 5.777The strategies developed in this section can be extended to the cases with $S\geq 3$ in principle, but with considerably more complex notations.
@@ 3.2 Identification by Disaggregated Data
where $\psi_{0,sr}=E[x_{s,t}|\varepsilon_{r,t}=1]-E[x_{s,t}|\varepsilon_{r,t}=0]$ and $\{\cdots\}$ is a linear combination of the elements in the curly brackets. We assume the unit effect normalization, $\psi_{0,ss}=1$ for $s=1,2$, so that a unit change in a shock results in a one-unit increase in its corresponding disaggregated variable. Since $x_{t}=x_{1,t}+x_{2,t}$, the baseline model (5) is a reduced SVMA model. The LP-IV estimand under the augmented SVMA model is
@@
for $s=1,2$ and $r\neq s$, provided that $\psi_{0,21}\psi_{0,12}\neq 1$. By imposing restrictions on $\psi_{0,sr}$, the coefficients $\alpha_{s}$ are identified. For the second IV, we impose the same restriction on $\psi_{0,sr}$ and identify the corresponding instrument-specific coefficients $\alpha_{s}$. Combining (16) across the two IVs then identifies $\theta_{h,y1}$ and $\theta_{h,y2}$.
@@ 3.3 Identification Bounds by Sign Restrictions
Panels (d)-(i) show the identified sets when the different sign restrictions, varying across the instruments, are imposed. In these cases multiple IV’s provide more informative bounds for the component-wise impulse responses without the inequality relationship between $\theta_{h,y1}/\theta_{0,x1}$ and $\theta_{h,y2}/\theta_{0,x2}$. Furthermore, the intersection of two identified sets with opposing sign restrictions gives the identified set bounded both above and below. This can happen when $\Theta_{+-}^{A}$ and $\Theta_{-+}^{B}$ intersect. If $\beta_{h}^{A}>\beta_{h}^{B}$, then the identified set for $(\theta_{h,y1}/\theta_{0,x1},\theta_{h,y2}/\theta_{0,x2})$ is given by
@@
Suppose that random variables $z_{t}^{A}$ and $z_{t}^{B}$ satisfy Assumption 1. We assume that $\text{cov}(z_{t}^{A},x_{t})>0$, $\text{cov}(z_{t}^{B},x_{t})>0$, and $\beta_{h}^{A}>\beta_{h}^{B}$ without loss of generality. [^8] If the covariance is not positive, we use $z_{t}$ as an instrument. Note that this will change the sign of $\alpha_{1}$ and $\alpha_{2}$ but not the sign of $\beta_{h}$.
@@ 3.4 Sharpness of Sign-Restricted LP-IV Identification
While we state sharp identification results for a bivariate VMA(1)(1) model, this is solely to simplify the exposition; our results hold for a more general VMA model, as discussed in the Supplemental Appendix.
@@ 3.4.1 Model Setup
Consider the SVMA(1)(1) model for the pair $(\mathbf{Y}_{t}^{\prime},\mathbf{Z}_{t}^{\prime})^{\prime}$ of $2\times 1$ endogenous and instrumental variable vectors $\mathbf{Y}_{t}=(x_{t},y_{t})^{\prime}$ and $\mathbf{Z}_{t}=(z_{t}^{A},z_{t}^{B})^{\prime}$ with respect to the $3\times 1$ structural shocks $\bm{\varepsilon}_{t}$ and the $2\times 1$ projection error vector $\mathbf{v}_{t}$:
@@
Note that this specification imposes the exclusion restriction that $\alpha_{3}^{j}=0$ for $j=A,B$, and $\alpha_{s}^{j}=d_{s}^{j}\lambda_{s}$ for $j=A,B$ and $s=1,2$. The parameters of interest are $\theta_{1,y1}/\theta_{0,x1}$ and $\theta_{1,y2}/\theta_{0,x2}$, the relative impulse responses of $y_{t}$ with respect to $\varepsilon_{1,t}$ and $\varepsilon_{2,t}$ at horizon $1$. We may consider the impulse responses at horizon 0 instead, and the results in Section 3.4.2 remain valid.
@@
[…]